OpenStudy (anonymous):

What is the additive inverse of -7x^2 + y? A. 7x^2 + y b. 7x^2 − y C. 7x − y d. -7x^2 − y

OpenStudy (larseighner):

Additive inverse of something is what when added to the something give 0 as a result. That is all you need to know. Example 3 additive inverse is -3 since 3 + (-3) = 0

OpenStudy (anonymous):

so its A?

OpenStudy (larseighner):

Test to see if you are right: -7x^2 + y + (7x^2 + y) = 2y But that is not zero. So, no.

OpenStudy (anonymous):

no..try otther option ...

OpenStudy (anonymous):

oh its B xD right?

OpenStudy (anonymous):

math is a subject that can be learn by practice not by just selecting...

OpenStudy (anonymous):

but im pretty sure its B ./..

OpenStudy (anonymous):

but i just want to be sure.

OpenStudy (anonymous):

if you put D then see it.. -7x^2-7x^2+y=-14x^2+y

OpenStudy (anonymous):

yuppp..its B...try it here..!!!

OpenStudy (larseighner):

Test to see if you are right -7x^2 + y +(7x^2 − y) = 0 So, yes.

OpenStudy (anonymous):

thanks ;d